40 students. 8 are left-handed. If 2 are chosen randomly what are the probabilities of the following events?
a. Both right-handed
b. Niether right-handed
c. One RH other LH
d. At least one is right-handed
Thanks a lot!

Answer by Lenzky(18)

N=40
Left handed =8
right handed =32
Use combination in your scientific calculator
a. 32C2 = 496
b. 8C2 = 28
c. 32C1x8C1 = 256
d. (32C1x8C1)+32C2 = 256+496 =752